# The electronic and magnetic structures of bilayer La3Ni2O7 at ambient pressure

### v2

@@ Abstract
We carry out a systematic study of the electronic and magnetic structure of the ambient-pressure bilayer La3Ni2O7. Employing the hybrid exchange-correlation functional, we show that the exchange-correlation pushes the antibonding $d_{z^{2}}$ bands below the Fermi level to be fully occupied. The calculated Fermi surfaces and the correlation normalized band structure match well with the experimental findings at ambient pressure. Moreover, the electronic susceptibility calculated for this new band structure features nesting-induced peaks near the wave vector $Q=(\pi/2,\pi/2)$, suggesting a possible density wave instability in agreement with recent experiments. Through a mean field study and DFT+U calculation, we confirm the spin-charge intertwined double stripe order is the magnetic ground state. Our results provide a faithful description for the low-pressure La3Ni2O7 electronic structure.
@@
Being a counterpart to copper, there has long been a theoretical proposition that nickelates could harbor high-temperature superconductivity (SC) rice_PhysRevB.59.7901 ; pickett_PhysRevB.70.165109 ; Khaliullin_PhysRevLett.100.016404 ; hujp_PhysRevX.5.041012 ; jphugene2016 , mirroring the physics observed in the high-$T_{c}$ cuprates doping_mott ; keimer_review ; sc_wtcu ; iron1 ; iron2 ; iron_review . However, identifying a suitable nickelate proves to be a complex challenge. Owing to sustained synthesis efforts nickelates0 ; nickelates1 ; nickelates2 ; nickelates4 ; nickelates5 , the manifestation of superconductivity has been successfully accomplished in thin films of the “infinite-layer” nickelates (Sr,Nd)NiO2 in 2019 lidanfeng ; lidanfeng2 ; lidanfeng3 , opening the Nickel age of superconductivity norman .
@@
In 2023, single crystals of a new type of bulk transition metal oxide La3Ni2O7 (LNO), structurally closer to hole-doped cuprates, were successfully synthesized meng_wang ; meng_wang2 ; chengjg_crystal . Under high pressure, a high-temperature superconducting transition $T_{c}\sim 80$K was reported meng_wang ; chengjg ; yuanhq ; chengjg_crystal ; chengjg_poly ; sunll , adding a new family to the nickelates superconductors. Following its discovery, extensive theoretical studies have been dedicated to understanding this novel material yaodx ; dagotto1 ; wangqh ; Kuroki ; guyh ; zhanggm ; werner ; yangf ; wucj ; dagotto2 ; yangyf2 ; ryee2023critical . However, the electronic structure of LNO and its SC origin are still under debate. Recently, a spin density wave order near $Q=(\pi/2,\pi/2)$ at ambient pressure is confirmed from resonant inelastic X-ray scattering (RIXS) rixs , muon spin rotation ($\mu$SR) musr ; musr2 , and nuclear magnetic resonance (NMR) nmr experiments below 153K meng_wang2 . Hence, the low-pressure electronic structure and magnetic structures are highly unexplored.
@@
In this work, we carry out a systematic study of the electronic and magnetic structure of LNO. We find that the crystal field splitting between the two $e_{g}$ orbitals and the bilayer coupling were highly underestimated in the previous DFT calculations. The important exchange-correlation effects turn out to be much stronger, resulting in large splittings that push the bonding $d_{z^{2}}$ bands below the Fermi level to become fully occupied. This leads to distinct new electronic structures for LNO. The calculated Fermi surfaces and the correlation normalized band structure match well with ARPES measurements at ambient pressure. In contrast, previous DFT results have not described this very well. Based on our new results, we find that LNO at ambient pressure harbors the density wave order close to the wave vector $Q$. Then, using the mean-field decoupling and DFT+U methods, we successfully identify the double stripe magnetic structure of LNO at ambient pressure.
@@
Much like the cuprates, the essential part of the LNO superconductor is the bilayer NiO2 block meng_wang , as illustrated in Fig. 1(a). We label them as the top (t) and bottom (b) layers. Around each Ni site, six oxygen atoms form a standard NiO6 octahedron. The two nearest neighbor octahedrons between the two layers are corner-shared by one apical oxygen. The LNO at ambient pressure is in the $Amam$ phase with the two octahedrons tilted in an alternating fashion, leading to four Ni atoms inside each unit cell. We label them as tA, tB, bA, bB with t/b indicating the layer index and A/B the sublattice index. This low-pressure (LP) phase evolves into the high symmetry $Fmmm$ structure phase under high pressure (HP). The two octahedrons line up with two Ni atoms per unit cell as shown in Fig. 1(a) and high-$T_{c}$ superconductivity emerges.
@@
The first-principle DFT calculation and its extensions are standard ways for determining the electronic structures and the crystal field energies. Various DFT calculations have been applied to finding the electronic structure of LNO meng_wang ; yaodx ; guyh ; dagotto1 , even long before the discovery of superconductivity in the HP phase Pickett . These calculations show that the bonding $d_{z^{2}}$ bands are active and cross the Fermi level ($E_{F}$) while the antibonding $d_{z^{2}}$ bands are empty, indicating an important role of $d_{z^{2}}$ orbitals. On the contrary, recent ARPES measurements find the bonding $d_{z^{2}}$ bands are below $E_{F}$ zhouxj , which deviates from previous calculations. We notice that most of the above-mentioned DFT calculations use the Generalized Gradient Approximation (GGA) exchange-correlation functional or its Perdew-Burke-Ernzerhof (PBE) version perdew1996generalized . However, it is widely known that the suitable exchange-correlation functional $E_{xc}$ is crucial for determining band gaps and binding energies martin2020electronic . For example, the GGA functional systematically underestimates band gaps of semiconductors martin2020electronic ; exchange_correlation . Because an accurate determination of the positions of the $d_{z^{2}}$ bands is crucial for understanding the electronic structure and correlated electronic states at ambient pressure, we apply a more accurate exchange-correlation functional to determine the interlayer binding and crystal splitting in LNO from the DFT perspective.
@@
Our DFT calculations employ the Vienna ab-initio simulation package (VASP) code kresse1996efficient with the projector augmented wave (PAW) method kresse1999ultrasoft . The hybrid functionals that combine the Hartree-Fock (HF) and Kohn-Sham (KS) theories becke1993density are used here. Typically, this can be more accurate compared to the semilocal GGA method exchange_correlation ; Jacob_ladder . Here, we implement the widely used HSE06 hybrid functional krukau2006influence . We also tested other unscreened hybrid functionals such as PBE0 perdew1996rationale ; ernzerhof1999assessment ; adamo1999toward and B3LYP stephens1994ab . As discussed in the supplemental material, all the hybrid functionals lead to similar results. The cutoff energy for expanding the wave functions into a plane-wave basis is set to be 500 eV. The energy convergence criterion is 10-6 eV. All calculations are conducted using the primitive cell to save time. The $\Gamma$-centered 5$\times$5$\times$5 k-meshes are used. The band structures for the LP phase using the HSE06 functional is plotted in Figs. 1(d). Compared to the GGA results (Fig. S1), the bonding $d_{z^{2}}$ bands are significantly pushed below $E_{F}$ by about $505$ meV ($\Gamma$ point in Fig. 1(d)).
@@
Figure 2: (a) The DFT band structure based on HSE06 hybrid functiona at ambient pressure (blue lines) and the corresponding TB bands (red lines). (b) The FSs at ambient pressure from the TB bands in (a). The red circles are experimental FSs from laser-based ARPES, while the green squares are data from synchrotron-based ARPES zhouxj . Notice that the $\Gamma$ electron pocket nearly nests with the $X_{1}$ hole pocket at $Q=(\pi/2,\pi/2)$. (c) The spectral function $A(k,\omega)$ calculated using DMFT at $U$ = $5.5$ eV, $J_{H}$ = $0.65$ eV, and $\beta=50$ based on TB bands in (a). The red crosses in (c) are experimental data from ARPES measurements zhouxj . (d) The spin susceptibility $\chi(q)$ is plotted as a function of $q$, which peaks around $Q=(\pi/2,\pi/2)$ ($X_{1}$ or $X_{2}$).
@@
Focusing on the partially occupied $e_{g}$ orbitals from the DFT calculations, we use the Wannier90 code mostofi2008wannier90 ; marzari2012maximally to fit the $e_{g}$ bands and extract the tight-binding (TB) model parameters. In the following discussion, we label the $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbital as $d_{x}$ and $d_{z}$ respectively. Due to the different octahedron environments, we define every $d_{x/z}$ orbital in its local octahedra coordinates rather than the global coordinates. Because of the 4 atoms per unit cell, we obtain an 8-band TB model, whose Hamiltonian $H_{LP}$ and hopping parameters are given in the supplemental material (Tab. S1). The dispersion of the TB $e_{g}$ bands is compared to the DFT calculations in Fig. 2(a), which confirms that the TB model provides a faithful description of the low-energy electron structure.
@@
The Fermi surfaces (FSs) of the TB model in the LP phase are plotted in Fig. 2(b). There are four FSs, one electron pocket centered around $\Gamma$, another small electron pocket around $X$, and two hole pockets centering around $X_{1}$ and $X_{2}$, respectively. To directly compare to experimental findings, we also plot the FS data obtained by angle-resolved photoemission spectroscopy (ARPES) recently in Fig. 2(b) zhouxj . The red circles are data from laser-based ARPES and the green squares are data from synchrotron-based ARPES. Remarkably, our theoretical FSs obtained by DFT and HSE06 exchange-correlation functional show excellent agreement with the ARPES data.
@@
To further the understanding of the electronic structure in the LP phase, we attempt a theoretical description of the band dispersions measured by ARPES zhouxj . To this end, it is necessary to include the electronic correlations beyond DFT, since the latter are known to produce the important renormalization of the band structure. We thus consider the multi-orbital Hubbard interactions,
@@
where $l=t,b$ is the layer index, $i$ is the site index in each layer and $\eta=x,z$ is the orbital index. To obtain the correlation-induced band renormalization, we apply the dynamical mean field theory (DMFT) to the full Hamiltonian $H_{LP}+H_{I}$. Specifically, we implement the DMFT calculation using the open-source TRIQS triqs package and its continuous-time quantum Monte Carlo as an impurity solver ctqmc . The obtained DMFT results at $U=5.5$eV and $J_{H}=0.65$eV are plotted in Fig. 2(c) along the $X-\Gamma-X$ direction. The ARPES data zhouxj are superimposed as red crosses in Fig. 2(c), which are very well described by the DMFT results. Despite the significant correlation-induced band renormalization necessary for the remarkable agreement with ARPES, we find that correlations do not cause appreciable inter-band carrier transfer and the interacting FSs are slightly changed from Fig. 2(b). Therefore, DMFT can qualitatively capture the band renormalization effects. This overall agreement with the experimental findings reassures that our new DFT electronic structure captured by the TB models plus the multi-orbital Hubbard interactions provide a more reliable description for the physics of La3Ni2O7 in its LP phase.
However, it is also important to emphasize that the energy bands calculated using the traditional GGA functional do not match well with the ARPES results even after considering correlation effects. Previously, many attempts using the DFT+U calculations based on the GGA functional have applied to the LP phase zhouxj . Although DFT+U can push the bands below the Fermi level, it fails to capture the band renormalization effect as shown in Fig. S3(b). We also apply the DMFT calculation with the TB inputs from traditional GGA functional. Although this method can capture the band renormalization effect, it fails to push the bands below the Fermi level as shown in Fig. S3(a).
Surprisingly, we find that the $\Gamma$ electron FS nearly nests with the $X_{1}$ hole FS, as marked in Fig. 2(b). The nesting between electron and hole FS pockets implies divergent electronic susceptibility at the nesting wave vector and potential symmetry-breaking density wave order at low temperatures. The calculated spin susceptibility $\chi(q)$ using TB bands from Fig. 2(a) is shown in Fig. 2(d) at different wave vectors, which indeed peaks around the near-nesting wave vector $Q=(\pi/2,\pi/2)$. This intriguing spin susceptibility feature is consistent with experimental findings rixs ; musr ; musr2 ; nmr .
@@
On the other hand, the above mean field cannot fully take care of the charge channel, which needs to be feedback to the electronic structure. A better way is using the simplified rotation invariant approach to the DFT+U, introduced by Dudarev $et$ $al.$ dudarev1998electron . Our calculations here are still based on the HSE06 hybrid functional. We also consider the above three stripe states and G-type. We find that within the range of $U=0$ to $U=4$, only the double-stripe configuration remains stable. The other two configurations spontaneously converge to the double stripe order. Therefore, we can be convinced that the double-stripe configuration is the magnetic ground state. Interestingly, the double-stripe order found here is a spin-charge intertwined state. The Ni atoms in the double-stripe order are divided into two groups with different magnetic moments and different charge densities. The larger charge atoms host relatively larger magnetic moments, as illustrated in Fig. 3(a). Therefore, a charge density wave order at $(\pi,\pi)$ emerges from the spin density wave ordering at $Q=(\pi/2,\pi/2)$.
@@
In summary, we presented a systematic study on the correlated electronic structure and magnetic properties in LP bilayer La3Ni2O7. Utilizing advanced exchange-correlation functionals, new electronic structures of LNO have been identified from the DFT calculations for the LP phases. In sharp contrast to previous GGA functional results, the bilayer coupling and the $e_{g}$ crystal fields are greatly enhanced. This enhancement pushes the bonding $d_{z^{2}}$ band down to lie below the Fermi level. The obtained Fermi surface structure in the LP phase matches well the recent ARPES measurements performed at ambient pressure. Beyond the DFT calculation, we apply the DMFT to study the effects of multi-orbital Hubbard interactions using the DFT (HSE06) derived TB models in the low-pressure phase. The renormalized band structures quantitatively match the bands detected by the ARPES-measured spectral functions. Moreover, using the new TB model, the calculated spin susceptibility exhibits near nesting peak structures around the wave vector $Q=(\pi/2,\pi/2)$, which is consistent with recent RIXS resolved magnon dispersion rixs and additional evidence from the $\mu$SR and NMR experiments musr ; musr2 ; nmr . Furthermore, using the DFT (HSE06)+U method and mean-field decoupling, we confirm that the ground state of the magnetic structure should be the double stripe order. A charge density wave intertwined with this SDW is also found, which calls for experimental verification. Therefore, we think our new TB model plus multi-orbital Hubbard interactions can provide a more faithful description of the electronic and magnetic structure in LNO.

### v3

@@ Abstract
We carry out a systematic study of the electronic and magnetic structure of the ambient-pressure bilayer La3Ni2O7. Employing the hybrid exchange-correlation functional, we show that the exchange-correlation pushes the bonding $d_{z^{2}}$ bands below the Fermi level to be fully occupied. The calculated Fermi surfaces and the correlation normalized band structure match well with the experimental findings at ambient pressure. Moreover, the electronic susceptibility calculated for this new band structure features nesting-induced peaks near the wave vector $Q=(\pi/2,\pi/2)$, suggesting a possible density wave instability in agreement with recent experiments. Through a mean field study and DFT+U calculation introducing a Hubbard U interaction within conventional density functional theory, we confirm the spin-charge intertwined double stripe order is the magnetic ground state. Our results provide a faithful description for the low-pressure La3Ni2O7 electronic structure.
I INTRODUCTION
@@
Being a counterpart to copper, there has long been a theoretical proposition that nickelates could harbor high-temperature superconductivity (SC) Anisimov et al. (1999); Lee and Pickett (2004); Chaloupka and Khaliullin (2008); Hu et al. (2015); Hu (2016), mirroring the physics observed in the high-$T_{c}$ cuprates Lee et al. (2006); Keimer et al. (2015); Maeno et al. (1994); Kamihara et al. (2008); Takahashi et al. (2008); Paglione and Greene (2010). However, identifying a suitable nickelate proves to be a complex challenge. Owing to sustained synthesis efforts Crespin et al. (1983); Greenblatt (1997); Hayward et al. (1999); Crespin et al. (2005); Kawai et al. (2010), the manifestation of superconductivity has been successfully accomplished in thin films of the “infinite-layer” nickelates (Sr,Nd)NiO2 in 2019 Li et al. (2019); Osada et al. (2020); Li et al. (2020), opening the Nickel age of superconductivity Norman (2020).
@@
In 2023, single crystals of a new type of bulk transition metal oxide La3Ni2O7 (LNO), structurally closer to hole-doped cuprates, were successfully synthesized Sun et al. (2023); Liu et al. (2022); Wang et al. (2023). Under high pressure, a high-temperature superconducting transition $T_{c}\sim 80$K was reported Sun et al. (2023); Hou et al. (2023); Zhang et al. (2024); Wang et al. (2023, 2024); Zhou et al. , adding a new family to the nickelates superconductors. Following its discovery, extensive theoretical studies have been dedicated to understanding this novel material Luo et al. (2023); Zhang et al. (2023); Yang et al. (2023a); Sakakibara et al. (2024); Gu et al. (2023); Yang et al. (2023b); Christiansson et al. (2023); Liu et al. (2023); Lu et al. (2024); Zhang et al. (2024); Yang et al. (2023c); Ryee et al. (2023); Jiang et al. (2024); Lechermann et al. (2023); Ouyang et al. (2024). However, the electronic structure of LNO and its SC origin are still under debate. Recently, a spin density wave order near $Q=(\pi/2,\pi/2)$ at ambient pressure is confirmed from resonant inelastic X-ray scattering (RIXS) Chen et al. (2024), muon spin rotation ($\mu$SR) Chen et al. (2024); Khasanov et al. (2024), and nuclear magnetic resonance (NMR) Dan et al. (2024) experiments below 153K Liu et al. (2022). Hence, the low-pressure electronic structure and magnetic structures are highly unexplored.
@@
In this work, we carry out a systematic study of the electronic and magnetic structure of LNO. We find that the crystal field splitting between the two $e_{g}$ orbitals and the bilayer coupling were highly underestimated in the previous density functional theory (DFT) calculations. The important exchange-correlation effects turn out to be much stronger, resulting in large splittings that push the bonding $d_{z^{2}}$ bands below the Fermi level to become fully occupied. This leads to distinct new electronic structures for LNO. The calculated Fermi surfaces and the correlation normalized band structure match well with ARPES measurements at ambient pressure. In contrast, previous DFT results have not described this very well. Based on our new results, we find that LNO at ambient pressure harbors the density wave order close to the wave vector $Q$. Then, using the mean-field decoupling and DFT+U methods, we successfully identify the double stripe magnetic structure of LNO at ambient pressure.
@@
II CRYSTAL STRUCTURE
Much like the cuprates, the essential part of the LNO superconductor is the bilayer NiO2 block Sun et al. (2023), as illustrated in Fig. 1(a). We label them as the top (t) and bottom (b) layers. Around each Ni site, six oxygen atoms form a standard NiO6 octahedron. The two nearest neighbor octahedrons between the two layers are corner-shared by one apical oxygen. The LNO at ambient pressure is in the $Amam$ phase with the two octahedrons tilted in an alternating fashion, leading to four Ni atoms inside each unit cell. We label them as tA, tB, bA, bB with t/b indicating the layer index and A/B the sublattice index. This low-pressure (LP) phase evolves into the high symmetry $Fmmm$ structure phase under high pressure (HP). The two octahedrons line up with two Ni atoms per unit cell as shown in Fig. 1(a) and high-$T_{c}$ superconductivity emerges.
@@ III DFT RESULTS
The first-principles DFT calculation and its extensions are standard ways for determining the electronic structures and the crystal field energies. Various DFT calculations have been applied to finding the electronic structure of LNO Sun et al. (2023); Luo et al. (2023); Gu et al. (2023); Zhang et al. (2023), even long before the discovery of superconductivity in the HP phase Pardo and Pickett (2011). These calculations show that the bonding $d_{z^{2}}$ bands are active and cross the Fermi level ($E_{F}$) while the antibonding $d_{z^{2}}$ bands are empty, indicating an important role of $d_{z^{2}}$ orbitals. On the contrary, recent ARPES measurements find the bonding $d_{z^{2}}$ bands are below $E_{F}$ Yang et al. (2024), which deviates from previous calculations. We notice that most of the above-mentioned DFT calculations use the Generalized Gradient Approximation (GGA) exchange-correlation functional or its Perdew-Burke-Ernzerhof (PBE) version Perdew et al. (1996a). However, it is widely known that the suitable exchange-correlation functional $E_{xc}$ is crucial for determining band gaps and binding energies Martin (2020). For example, the GGA functional systematically underestimates band gaps of semiconductors Martin (2020); Borlido et al. (2019). Because an accurate determination of the positions of the $d_{z^{2}}$ bands is crucial for understanding the electronic structure and correlated electronic states at ambient pressure, we apply a more accurate exchange-correlation functional to determine the interlayer binding and crystal splitting in LNO from the DFT perspective.
@@
Our DFT calculations employ the Vienna ab-initio simulation package (VASP) code Kresse and Furthmüller (1996) with the projector augmented wave (PAW) method Kresse and Joubert (1999). The hybrid functionals that combine the Hartree-Fock (HF) and Kohn-Sham (KS) theories Becke (1993) are used here. Typically, this can be more accurate compared to the semilocal GGA method Borlido et al. (2019); Perdew and Schmidt (2001). Here, we implement the widely used HSE06 hybrid functional Krukau et al. (2006). We also tested other unscreened hybrid functionals such as PBE0 Perdew et al. (1996b); Ernzerhof and Scuseria (1999); Adamo and Barone (1999) and B3LYP Stephens et al. (1994). As discussed in the supplemental material (SM) SM , all the hybrid functionals lead to similar results. The cutoff energy for expanding the wave functions into a plane-wave basis is set to be 500 eV. The energy convergence criterion is 10-6 eV. All calculations are conducted using the primitive cell to save time. The $\Gamma$-centered 5$\times$5$\times$5 k-meshes are used. The band structures for the LP phase using the HSE06 functional is plotted in Figs. 1(d). Compared to the GGA results (See Fig. S1 in SM SM ), the bonding $d_{z^{2}}$ bands are significantly pushed below $E_{F}$ by about $505$ meV ($\Gamma$ point in Fig. 1(d)).
@@
Figure 2: (a) The DFT band structure based on HSE06 hybrid functiona at ambient pressure (blue lines) and the corresponding TB bands (red lines). (b) The FSs at ambient pressure from the TB bands in (a). The red circles are experimental FSs from laser-based ARPES, while the green squares are data from synchrotron-based ARPES Yang et al. (2024). Notice that the $\Gamma$ electron pocket nearly nests with the $X_{1}$ hole pocket at $Q=(\pi/2,\pi/2)$. (c) The spectral function $A(k,\omega)$ calculated using DMFT at $U$ = $5.5$ eV, $J_{H}$ = $0.65$ eV, and $\beta=50$ based on TB bands in (a). The red crosses in (c) are experimental data from ARPES measurements Yang et al. (2024). (d) The real part of the electron self-energy from DMFT calculation. (e) The negative imaginary part of the electron self-energy from DMFT calculation. (f) The spin susceptibility $\chi(q)$ is plotted as a function of $q$, which peaks around $Q=(\pi/2,\pi/2)$ ($X_{1}$ or $X_{2}$).
@@
Focusing on the partially occupied $e_{g}$ orbitals from the DFT calculations, we use the Wannier90 code Mostofi et al. (2008); Marzari et al. (2012) to fit the $e_{g}$ bands and extract the tight-binding (TB) model parameters. In the following discussion, we label the $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbital as $d_{x}$ and $d_{z}$ respectively. Due to the different octahedron environments, we define every $d_{x/z}$ orbital in its local octahedra coordinates rather than the global coordinates. Because of the 4 atoms per unit cell, we obtain an 8-band TB model, whose Hamiltonian $H_{LP}$ and hopping parameters are given in the Tab. S1 in SM SM . The dispersion of the TB $e_{g}$ bands is compared to the DFT calculations in Fig. 2(a), which confirms that the TB model provides a faithful description of the low-energy electron structure.
@@
The Fermi surfaces (FSs) of the TB model in the LP phase are plotted in Fig. 2(b). There are four FSs, one electron pocket centered around $\Gamma$, another small electron pocket around $X$, and two hole pockets centering around $X_{1}$ and $X_{2}$, respectively. To directly compare to experimental findings, we also plot the FS data obtained by angle-resolved photoemission spectroscopy (ARPES) recently in Fig. 2(b) Yang et al. (2024). The red circles are data from laser-based ARPES and the green squares are data from synchrotron-based ARPES. Remarkably, our theoretical FSs obtained by DFT and HSE06 exchange-correlation functional show excellent agreement with the ARPES data.
IV DMFT RESULTS
@@
To further the understanding of the electronic structure in the LP phase, we attempt a theoretical description of the band dispersions measured by ARPES Yang et al. (2024). To this end, it is necessary to include the electronic correlations beyond DFT, since the latter are known to produce the important renormalization of the band structure. We thus consider the multi-orbital Hubbard interactions,
@@
where $l=t,b$ is the layer index, $i$ is the site index in each layer and $\eta=x,z$ is the orbital index. To obtain the correlation-induced band renormalization, we apply the dynamical mean field theory (DMFT) to the full Hamiltonian $H_{LP}+H_{I}$. Specifically, we implement the DMFT calculation using the open-source TRIQS Parcollet et al. (2015) package and its continuous-time quantum Monte Carlo as an impurity solver Seth et al. (2016). The obtained DMFT results at $U=5.5$eV and $J_{H}=0.65$eV are plotted in Fig. 2(c) along the $X-\Gamma-X$ direction. The ARPES data Yang et al. (2024) are superimposed as red crosses in Fig. 2(c), which are very well described by the DMFT results. Despite the significant correlation-induced band renormalization necessary for the remarkable agreement with ARPES, we find that correlations do not cause appreciable inter-band carrier transfer and the interacting FSs are slightly changed from Fig. 2(b).
[…]
$\omega-\epsilon=\rm{Re}\Sigma(\omega).$ (2)
[…]
These overall agreements with the experimental findings reassures that our new DFT electronic structure captured by the TB models plus the multi-orbital Hubbard interactions provide a reliable description for the physics of La3Ni2O7 in its LP phase.
However, it is also important to emphasize that the energy bands calculated using the traditional GGA functional do not match well with the ARPES results even after considering correlation effects. Previously, many attempts using the DFT+U calculations based on the GGA functional have applied to the LP phase Yang et al. (2024). Although DFT+U can push the bands below the Fermi level, it fails to capture the band renormalization effect as shown in Fig. S3(b) in SM SM . We also apply the DMFT calculation with the TB inputs from traditional GGA functional. Although this method can capture the band renormalization effect, it fails to push the bands below the Fermi level as shown in Fig. S3(a). The more detailed information can also be seen in SM SM .
V MAGNETIC STRUCTURE
Surprisingly, we find that the $\Gamma$ electron FS nearly nests with the $X_{1}$ hole FS, as marked in Fig. 2(b). The nesting between electron and hole FS pockets implies divergent electronic susceptibility at the nesting wave vector and potential symmetry-breaking density wave order at low temperatures. The calculated spin susceptibility $\chi(q)$ using TB bands from Fig. 2(a) is shown in Fig. 2(f) at different wave vectors, which indeed peaks around the near-nesting wave vector $Q=(\pi/2,\pi/2)$. This intriguing spin susceptibility feature is consistent with experimental findings Chen et al. (2024, 2024); Khasanov et al. (2024); Dan et al. (2024).
@@ V MAGNETIC STRUCTURE
On the other hand, the above mean field cannot fully take care of the charge channel, which needs to be feedback to the electronic structure. A better way is using the simplified rotation invariant approach to the DFT+U, introduced by Dudarev $et$ $al.$ Dudarev et al. (1998). Our calculations here are still based on the HSE06 hybrid functional. We also consider the above three stripe states and G-type. We find that within the range of $U=0$eV to $U=4$eV, only the double-stripe configuration remains stable. The other two configurations spontaneously converge to the double stripe order. Therefore, we can be convinced that the double-stripe configuration is the magnetic ground state. Interestingly, the double-stripe order found here is a spin-charge intertwined state. The Ni atoms in the double-stripe order are divided into two groups with different magnetic moments and different charge densities. The larger charge atoms host relatively larger magnetic moments, as illustrated in Fig. 3(a). Therefore, a charge density wave order at $(\pi,\pi)$ emerges from the spin density wave ordering at $Q=(\pi/2,\pi/2)$.
@@
VI SUMMARY
Although the main focus of our paper is the electronic structure and magnetism of the LP phase, we would like to briefly mention the understanding of superconductivity in the HP using the HSE06 functional. In previous studies Luo et al. (2023); Zhang et al. (2023); Yang et al. (2023a); Sakakibara et al. (2024); Gu et al. (2023); Yang et al. (2023b); Christiansson et al. (2023); Liu et al. (2023); Lu et al. (2024); Zhang et al. (2024); Yang et al. (2023c); Ryee et al. (2023); Jiang et al. (2024); Lechermann et al. (2023), it has been suggested that the bonding state of the $d_{z}$ orbital plays a crucial role in superconducting pairing, as it remains active at the Fermi surface. There is also a debate regarding the $s^{\pm}$-wave Yang et al. (2023a); Gu et al. (2023); Zhang et al. (2024) and $d$-wave pairing symmetries Jiang et al. (2024); Lechermann et al. (2023). However, our use of the HSE06 functional indicates that in the LP phase, the bonding state of $d_{z}$ is fully occupied, meaning it no longer appears at the Fermi surface. In fact, we also calculated the band structure of the HP phase using the same method (with the crystal structure obtained from Sun et al. (2023)), as shown in Fig. S5 in SM SM . Due to band unfolding, the band top of the bonding state of $d_{z}$ is located at the M point, and it is clearly still below the Fermi level. Therefore, we believe that based on our band structure, the superconducting order parameter is likely to exhibit $d$-wave symmetry, which is consistent with our previous theoretical analysis Jiang et al. (2024). Currently, there are first-principles techniques that, by combining DFT-calculated band structures with pairing assumptions, provide further insight Csire, Gábor and Újfalussy, Balázs and Annett, James F (2018). It is also worth further investigating the impact of band structure on the superconducting ground state using this approach.
In summary, we presented a systematic study on the correlated electronic structure and magnetic properties in LP bilayer La3Ni2O7. Utilizing advanced exchange-correlation functionals, new electronic structures of LNO have been identified from the DFT calculations for the LP phases. In sharp contrast to previous GGA functional results, the bilayer coupling and the $e_{g}$ crystal fields are greatly enhanced. This enhancement pushes the bonding $d_{z^{2}}$ band down to lie below the Fermi level. The obtained Fermi surface structure in the LP phase matches well the recent ARPES measurements performed at ambient pressure. Beyond the DFT calculation, we apply the DMFT to study the effects of multi-orbital Hubbard interactions using the DFT (HSE06) derived TB models in the low-pressure phase. The renormalized band structures quantitatively match the bands detected by the ARPES-measured spectral functions. Moreover, using the new TB model, the calculated spin susceptibility exhibits near nesting peak structures around the wave vector $Q=(\pi/2,\pi/2)$, which is consistent with recent RIXS resolved magnon dispersion Chen et al. (2024) and additional evidence from the $\mu$SR and NMR experiments Chen et al. (2024); Khasanov et al. (2024); Dan et al. (2024). Furthermore, using the DFT (HSE06)+U method and mean-field decoupling, we confirm that the ground state of the magnetic structure should be the double stripe order. A charge density wave intertwined with this SDW is also found, which calls for experimental verification. Therefore, we think our new TB model plus multi-orbital Hubbard interactions can provide a more faithful description of the electronic and magnetic structure in LNO.